MIRR- A project has an initial cost of $70,225, expected net cash inflows of $10,000 per year for seven years, and the cost of capital of 11%. What is the projects MIRR?
Answer #1

Project MIRR

The modified internal rate of return (MIRR) is a financial measure of an investment's attractiveness. It is used in capital budgeting to rank alternative investments of equal size. As the name implies, MIRR is a modification of the internal rate of return (IRR) and as such aims to resolve some problems with the IRR.

In the above case

MIRR=

outlay= terminal cash flow/(1+mirr)n

outlay(o)=70225

terminal cash flow(P)=10000

n=7 years

= (70225/70000)=1/(1+mirr)n×1.11

mirr=5.865